# Learning Heterogeneous Mixture of Scene Experts for Large-scale Neural Radiance Fields

### v1

@@ Abstract
Recent Neural Radiance Field (NeRF) methods on large-scale scenes have demonstrated promising results and underlined the importance of scene decomposition for scalable NeRFs. Although these methods achieved reasonable scalability, there are several critical problems remaining unexplored in the existing large-scale NeRF modeling methods, i.e., learnable decomposition, modeling scene heterogeneity, and modeling efficiency. In this paper, we introduce Switch-NeRF++, a Heterogeneous Mixture of Hash Experts (HMoHE) network that addresses these challenges within a unified framework. Our framework is a highly scalable NeRF that learns heterogeneous decomposition and heterogeneous Neural Radiance Fields efficiently for large-scale scenes in an end-to-end manner. In our framework, a gating network learns to decomposes scenes into partitions and allocates 3D points to specialized NeRF experts. This gating network is co-optimized with the experts, by our proposed Sparsely Gated Mixture of Experts (MoE) NeRF framework. Our network architecture incorporates a hash-based gating network and distinct heterogeneous hash experts. The hash-based gating efficiently learns the decomposition of the large-scale scene. The distinct heterogeneous hash experts consist of hash grids of different resolution ranges. This enables effective learning of the heterogeneous representation of different decomposed scene parts within large-scale complex scenes. These design choices make our framework an end-to-end and highly scalable NeRF solution for real-world large-scale scene modeling to achieve both quality and efficiency. We evaluate our accuracy and scalability on existing large-scale NeRF datasets. Additionally, we also introduce a new dataset with very large-scale scenes ($>6.5km^{2}$) from UrbanBIS. Extensive experiments demonstrate that our approach can be easily scaled to various large-scale scenes and achieve state-of-the-art scene rendering accuracy. Furthermore, our method exhibits significant efficiency gains, with an 8x acceleration in training and a 16x acceleration in rendering compared to the best-performing competitor Switch-NeRF. The codes and trained models will be released in https://github.com/MiZhenxing/Switch-NeRF.
@@ 1 Introduction
Neural Radiance Fields (NeRF) [1] have shown the powerfulness for high-fidelity 3D scene modeling. Recent methods [2, 3, 4] have scaled NeRF to large-scale scenes with thousands of high-resolution images. The scalability of them essentially comes from the decomposition of the 3D scenes. Several sub-NeRF-networks handle different partitions of a large scene. Mega-NeRF [2] and Block-NeRF [3] both incorporate hand-crafted decomposition rules for different scenes. They decompose 3D scenes by clustering based on 3D physical distances or street blocks.
Despite the promising performances achieved by current large-scale NeRF methods, several critical challenges remain unexplored. Firstly, design a universal decomposition rule is extremely challenging for different large-scale scenes. The real-world large-scale scenes tend to have very different and complex scene structures. Hand-crafted rules will inevitably brings adaptation issues for distinct scenarios. Moreover, the non-learnable nature of hand-crafted decomposition restricts the network’s ability to optimize the scene decomposition and radiance fields together within an end-to-end framework. To handle this problem, we need a learnable decomposition method to automatically learn different decomposition for different scenes and we can optimize the decomposition together with the scene representation. Secondly, a large-scale scene is typically irregular and complex, with different scene parts exhibiting diverse appearance and geometry distributions, which requires heterogeneous modeling capabilities to produce high-quality scene representation. However, current large-scale NeRF methods often employ identical networks, such as Multilayer Perceptrons (MLPs), for different parts of the scene, lacking explicit designs for heterogeneous modeling. Thirdly, the long training time and high GPU memory usage of existing methods is a severe bottleneck that limits their scalability when dealing with much larger scenes. Although several methods such as Instant-NGP [5] have been proposed to enhance the NeRF representation and accelerate the NeRF training on small- and large-scale scenes [6, 7, 8, 9], they still attempt to fit an entire scene into a single global network, overlooking the crucial aspect of heterogeneous scene representation from distinct scene decompositions. Consequently, these methods suffer from clear limitations in scalability and representation effectiveness, particularly when dealing with very large scenes.
@@
An sparse framework for unified optimization of scene decomposition and neural radiance fields. We introduce an end-to-end sparse framework that unifies optimization of scene decomposition and Neural Radiance Fields by a learnable gating mechanism. It employs a trainable gating network to dynamically select and sparsely activate specialized sub-networks for individual 3D points. The whole network is trained end-to-end without heuristic supervision and priors of the 3D scene shape or the distribution of scene images. This enables scalable and generic modeling of large-scale scenes. A critical challenge of training our network end-to-end lies in the discrete selection operation of sub-networks. This operation is not differentiable and the gradients cannot be back-propagated through it into the gating network. Inspired by Sparsely-Gated Mixture-of-Experts (MoE) [11], we address this problem by multiply gate values to the expert outputs, adding a connection between the gating network and experts. For the scene representation, distinct scene regions are handled by different NeRF experts (sub-networks). The gating network routes 3D points to relevant experts based on gate values. A unified output head is designed to unify the prediction of multiple NeRF experts. This head impose feature alignment among the implicit features of NeRF experts, resolving inconsistencies among different partitions.
Heterogeneous Mixture of Hash Experts (HMoHE) to model heterogeneous scene representations. Previous large-scale NeRF methods typically use the same network structure for different sub-networks. However, different parts of the large-scale scenes have heterogeneous and diverse distributions, which need heterogeneous representations. In our framework, the learned decomposition of the entire scene are represented by a mixture of heterogeneous hash experts (HMoHE). Each decomposed scene part is handled by a distinct expert network. To ensure the diversity and flexibility of representation, we design the hash experts as heterogeneous. We make the experts exploit diverse resolution ranges, allowing our network to select a suitable resolution range for each scene part.
@@
Effcient implement of hash-based gating and dispatching. To implement HMoHE more efficiently, we seamlessly integrate the dispatching of 3D points into the CUDA implementation of hash. We merge the hashing and dispatching of points together and avoid extra allocation of memories. This ensures that the point dispatching does not significantly impact the training efficiency.
@@
This paper significantly extends our previous ICLR version, i.e. Switch-NeRF [4]. Specifically, this work introduces a brand new design of Heterogeneous Mixture of Hash framework. We redesign the gating network and experts in Switch-NeRF exploiting highly efficient hash networks. We propose to learn heterogeneous decomposition and representation with the hash gating network and heterogeneous experts. For the point dispatching, we integrate it into the hash operation. We also adapt a new large-scale scene for better evaluation. Comparing to Switch-NeRF, our new framework effectively models the heterogeneous representation of large-scale scenes, and significantly improves the accuracy and efficiency. The proposed method achieves an 8x acceleration in training and a 16x acceleration in rendering compared to Switch-NeRF.
@@
  • We introduce a novel framework for large-scale neural radiance fields (NeRF). It integrates learnable decomposition, heterogeneous scene decomposition and representation, and efficient modeling in an end-to-end NeRF model, which is general applicable and does not need hand-crafted decomposition and composition rules.
@@
  • We design an efficient hash-based gating network to automatically learn efficient scene decomposition, and model scene partitions by the proposed Heterogeneous Mixture of Hash Experts (HMoHE). The hash-based gating and HMoHE are jointly learned to achieve heterogeneous and highly scalable NeRF representations of large-scale scenes.
@@
  • We present efficient implementation of the framework to make it a highly scalable and efficient learning framework that can handle extremely large-scale scenes. We adapt a new dataset for benchmarking highly scalable NeRF scene modeling. Our NeRF framework achieves significant improvement in terms of scalability, accuracy, and efficiency compared to the best-performing methods such as Switch-NeRF.
@@
Figure 2: Large-scale NeRFs with different decomposition methods. Dotted lines indicate non-differentiable operations, while solid lines represent differentiable operations that can back-propagate gradients in training. Mega-NeRF [2] divides scene by clustering image pixels by 3D distances. Block-NeRF [3] divide images sets according to street blocks, which is actually based on camera distributions. In both kinds of methods, the decomposition are hand-crafted heuristics rules and not trainable. They require separate training for their sub-networks. Switch-NeRF introduces a unified framework where scene decomposition is learned automatically via a gating network, enabling end-to-end training of the entire framework. However, Switch-NeRF designs homogeneous experts (MLPs) with the same structures, which learns homogeneous representation of the scene. In contrast, Switch-NeRF++ employs efficient hash network as the gate network and designs heterogeneous hash experts to learn heterogeneous decomposition and representation of a large-scale scene.
@@
Figure 3: The framework overview of Switch-NeRF and Switch-NeRF++ for large-scale scene modeling. A 3D scene point x is first fed into the gating network to obtain a feature encoding $f_{g}$. $f_{g}$ is then sent to a small MLP head to predict gating values. The Top-$1$ index $i$ of the gate values is used in 3D point dispatching that determines which hash expert should be selected for the input point. In this way, the entire scene can be decomposed into different partitions. Then x is dispatched into its corresponding expert to encode its distinct distribution into an expert encoding $E_{i}(\textbf{x})$. Then we multiply the gate value $G_{i}(\textbf{x})$ on $E_{i}(\textbf{x})$ and obtain the final expert output $\hat{E}_{i}(\textbf{x})$, which is further fed into a small MLP head to predict the point density $\sigma$ and point color c. The spherical harmonics encoding (i.e., SH) of direction d, and the appearance embedding (i.e., AE) are used for view-dependent prediction of color c. In Switch-NeRF, the gating network and the NeRF experts are all MLP-based networks, which learns homo. In contrast, in Switch-NeRF++, we leverage the efficient Hash-based Gating Network to dispatch 3D points. Furthermore, we design heterogeneous hash Heterogeneous Hash experts which have different scale ranges, to explicitly model the heterogeneous distribution of the large-scale scene.
@@ 2 Related Work
Neural Radiance Fields. Neural Radiance Fields method (NeRF) [1] are proposed to encode a 3D scene into an MLP network for novel view synthesis, with the multiple images and their cameras as input. The implicit framework of NeRF is very simple and effective, which have been widely applied to many 3D reconstruction tasks.
@@
Some NeRF variants apply NeRF on more challenging scenes. NeRF++ [28] extends NeRF to unbounded scenes by separating a scene into a foreground and a background, and applying the inverted sphere parameterization on the background. NeRF in the wild [29] attaches an appearance embedding to each image. It conditions the NeRF network with these appearance embeddings to model the appearance variation in each image in the wild environment. Mip-NeRF [30] and Mip-NeRF 360 [31] address sampling and aliasing of bounded and 360 degree scenes by designing a new scale-aware positional encoding. DSNeRF [32] uses sparse depth from colmap to regularize and speed up NeRF training. SPARF [33] and CVT-xRF [34] largely reduce the view number used in NeRF training and produce high-quality NeRF with sparse image inputs.
@@
Although the original NeRF targets novel view synthesis, it is a flexible 3D representation encoding 3D geometry. Several methods have focus on improving the geometry quality of NeRF for accurate 3D surface reconstruction [35, 36]. NeuS [35] represents a surface as a signed distance function (SDF) and learns a SDF field by volume rendering. Neuralangelo [36] recovers high-quality surface details by higher-order derivatives and coarse-to-fine optimization on the hash grids.
@@
NeRF methods typically need camera parameters as input. These cameras are usually predicted by SFM methods such as Colmap [37]. Several methods try to estimate camera parameters together with the neural radiance fields. NeRF– [38] treats camera parameters as trainable model parameters and jointly cameras in NeRF training. BARF [39] finds that use coarse-to-fine positional encoding in training helps the convergence of camera parameters.
@@
The training and inference speed of NeRF are critical in real-world application. Many methods have been proposed for fast training and real-time rendering of NeRF. PlenOctrees [40], FastNeRF [41] and Kilo-NeRF [42] mainly speed up inference of NeRF. There are also methods speed up both the training and rendering of NeRF, such as the DVGO [6], Plenoxels [43], Instant-NGP [5], TensoRF [8] and 3D Gaussian Splatting [44]. Instant-NGP uses a multi-resolution hash encoding to encode a scene into a set of hash maps by interpolating trainable hash features. TensoRF [8] represents a 3D radiance field using trainable 2D plane features and 1D feature vectors. Different from typical NeRF methods using volume rendering, 3D Gaussian Splatting (3DGS) [44] used 3D Gaussians to represent a 3D scene. It renders images by projection and $\alpha$-blending, which is very fast and support high-resolution real-time rendering. 3DGS needs a point cloud from SfM or random initialization to initialize 3D Gaussians.
@@
The original NeRF needs to be trained on each scene and cannot generalize across different scenes. Pixel-NeRF [45] proposes to include image features as an additional condition and trains the model on a set of different scenes. MVSNeRF [46], NerfingMVS and IBRNet [47] et. al. handle unsupervised Multi-view Stereo problem by combining the volume rendering and Multi-view Stereo together. These methods take single view or Multi-view image features as input and can generalize to different scenes.
@@
The flexible NeRF representation have been applied to 3D generation [48, 49, 50] and editing tasks [51, 52]. EG3D [48] simplify the NeRF network into triplane feature maps, which has regular shapes and is easy for generation. LRM [49] propose to use a large transformer to generate triplanes from a single image.
@@
The above mentioned reconstruction methods typically works on small scale scenes. Our method targets to scale NeRF methods to large-scale scenes and handle the problems of learnable decomposition, and heterogeneous, scalable and efficient scene modeling for large-scale NeRF.
@@
Large-scale Neural Radiance Fields. When applying on large-scale NeRF, a critical problem of NeRF is that handling a large data volume needs a scalable framework and more network parameters. Recent large-scale NeRF methods mainly focus on decompose a large-scale scene into several parts and handle each part by a sub-network. Mega-NeRF [2] and Block-NeRF [3] successfully scale NeRF to urban-scale scenes sparsely by scene decomposition. Although achieving impressive accuracy, their scene decomposition is mainly conducted by heuristic hand-crafted rules. Mega-NeRF [2] places several centroids in a scene and divides the 3D points in the scene by 3D distances to these centroids. Block-NeRF [3] divides the training images into several subsets according to street blocks. These two methods train the sub-networks separately. Therefore, they need composition strategies such as distance-based weights to merge the rendering results. Different from methods using handcrafted decomposition methods, Switch-NeRF [4] firstly proposes an large-scale NeRF framework that automatically learns scene decomposition together with the Neural Radiance Fields in an end-to-end manner. It designs a sparse network structure containing a gating network and several NeRF experts, base on the mechanism of Mixture of Experts (MoE) [11]. It requires no prior knowledge of scene structures and decomposition rules. Our method extends from Switch-NeRF with brand-new designs of the hash-based network, heterogeneous experts and efficient hash dispatching. Our method remarkably improving the scalability, efficiency and accuracy of large-scale NeRFs compared to best-performing methods.
@@
Some other methods [53, 9] also use plane features and hash encodings on large-scale scenes. However, they still encode a large-scale scene into a global network so their scalability is still limited. The Gaussian splatting [54] has inspired large-scale NeRFs such as the Hierarchical-3D-Gaussians [55], CityGaussian [56], DoGaussian [57], VastGaussian [58], PyGS [59], etc. due to its efficiency. The Hierarchical-3D-Gaussians [55] uses a divide-and-conquer method that trains very large scenes in independent chunks. It organizes the chunks into a hierarchy and provides an efficient Level-of-Detail (LOD) rendering. CityGaussian [56] and VastGaussian [58] also use a divide-and-conquer method to manually divides the scene into smaller cells. In this paper, our method focuses on scaling to large-scale scenes by learning efficient scene decomposition and representations. Our method explores the better use of the multi-scale nature of grids to adaptively model the multi-scale nature of 3D scenes.
@@
Mixture of Experts (MoE). The methodology of Mixture of Experts has a long history for scaling the model size [60, 61]. Recent applications of MoE in deeplearning mainly follow the design of [11]. It replaces the feed-forward networks in a language model by Mixture-of-Experts layers. The experts are activated sparsely by a Top-$k$ gating network. Each sample will only be dispatched to $k$ experts while $k$ is typically set as 2 and is much smaller that the total number of experts. In order to avoids imbalanced training and training collapse, it regularize gate values by an gate loss which can make the training of different experts more balanced. The MoE network is extensively applied for Natural Language Processing (NLP) [62, 12] and Vision [63, 64] tasks to scale the capacity and capability of neural networks. GShard [62] standardize the dispatching operations by Einops [65, 62]. It supports local dispatching and parallel gating by designing a simpler balancing loss and defining the capacity factor for experts. The value of $k$ in MoE gating directly influence the sparsity, efficiency and accuracy of MoE. Top-$2$ typically can get good accuracy while maintaining sparsity [11, 62]. Switch Transformer [12] has investigated into the training recipes of MoE in NLP tasks. It concludes that under carefully designed training setting, Top-$2$ operation in MoE is not necessary. Top-$1$ in MoE can already get high accuracy while largely decrease the cost of computation and communication for sample dispatching. Several MoE methods have investigated different gating mechanisms, such as Hash Routing [66], BASE [67] and Expert Choice Routing [68].

### v2

@@ Abstract
Recent Neural Radiance Field (NeRF) methods on large-scale scenes have demonstrated promising results and underlined the importance of scene decomposition for scalable NeRFs. Although these methods achieved reasonable scalability, there are several critical problems remaining unexplored in the existing large-scale NeRF modeling methods, i.e., learnable decomposition, modeling scene heterogeneity, and modeling efficiency. In this paper, we introduce Switch-NeRF++, a Heterogeneous Mixture of Hash Experts (HMoHE) network that addresses these challenges within a unified framework. Our framework is a highly scalable NeRF that learns heterogeneous decomposition and heterogeneous Neural Radiance Fields efficiently for large-scale scenes in an end-to-end manner. In our framework, a gating network learns to decompose scenes into partitions and allocates 3D points to specialized NeRF experts. This gating network is co-optimized with the experts by our proposed Sparsely Gated Mixture of Experts (MoE) NeRF framework. Our network architecture incorporates a hash-based gating network and distinct heterogeneous hash experts. The hash-based gating efficiently learns the decomposition of the large-scale scene. The distinct heterogeneous hash experts consist of hash grids of different resolution ranges. This enables effective learning of the heterogeneous representation of different decomposed scene parts within large-scale complex scenes. These design choices make our framework an end-to-end and highly scalable NeRF solution for real-world large-scale scene modeling to achieve both quality and efficiency. We evaluate our accuracy and scalability on existing large-scale NeRF datasets. Additionally, we also introduce a new dataset with very large-scale scenes ($6.5km^{2}$) from UrbanBIS. Extensive experiments demonstrate that our approach can be easily scaled to various large-scale scenes and achieve state-of-the-art scene rendering accuracy. Furthermore, our method exhibits significant efficiency gains, with an 8x acceleration in training and a 16x acceleration in rendering compared to the best-performing competitor Switch-NeRF. The codes and trained models will be released in https://github.com/MiZhenxing/Switch-NeRF.
@@ 1 Introduction
Neural Radiance Fields (NeRF) [1] have shown the powerfulness for high-fidelity 3D scene modeling. Recent methods [2, 3, 4] have scaled NeRF to large-scale scenes with thousands of high-resolution images. The scalability of them essentially comes from the decomposition of the 3D scenes. Several sub-networks handle different partitions of a large scene. Mega-NeRF [2] and Block-NeRF [3] both incorporate hand-crafted decomposition rules for different scenes. They decompose 3D scenes by clustering based on 3D physical distances or street blocks.
Despite the promising performances achieved by current large-scale NeRF methods, several critical challenges remain unexplored. Firstly, designing a universal decomposition rule is extremely challenging for different large-scale scenes. The real-world large-scale scenes tend to have very different and complex scene structures. Hand-crafted rules will inevitably bring adaptation issues for distinct scenarios. Moreover, the non-learnable nature of hand-crafted decomposition restricts the network’s ability to optimize the scene decomposition and radiance fields together within an end-to-end framework. To handle this problem, we need a learnable decomposition method to automatically learn different decompositions for different scenes, and we can optimize the decomposition together with the scene representation. Secondly, a large-scale scene is typically irregular and complex, with different scene parts exhibiting diverse appearance and geometry distributions, which requires heterogeneous modeling capabilities to produce high-quality scene representation. However, current large-scale NeRF methods often employ identical networks, such as Multilayer Perceptrons (MLPs), for different parts of the scene, lacking explicit designs for heterogeneous modeling. Thirdly, the long training time and high GPU memory usage of existing methods are a severe bottleneck that limits their scalability when dealing with much larger scenes. Although several methods such as Instant-NGP [5] have been proposed to enhance the NeRF representation and accelerate the NeRF training on small- and large-scale scenes [6, 7, 8, 9], they still attempt to fit an entire scene into a single global network, overlooking the crucial aspect of heterogeneous scene representation from distinct scene decompositions. Consequently, these methods suffer from clear limitations in scalability and representation effectiveness, particularly when dealing with very large scenes.
@@
A sparse framework for unified optimization of scene decomposition and neural radiance fields. We introduce an end-to-end sparse framework that unifies optimization of scene decomposition and Neural Radiance Fields by a learnable gating mechanism. It employs a trainable gating network to dynamically select and sparsely activate specialized experts for individual 3D points. The whole network is trained end-to-end without heuristic supervision and priors of the 3D scene shape or the distribution of scene images. This enables scalable and generic modeling of large-scale scenes. A critical challenge of training our network end-to-end lies in the discrete selection operation of experts. This operation is not differentiable, and the gradients cannot be back-propagated through it into the gating network. Inspired by Sparsely-Gated Mixture-of-Experts (MoE) [11], we address this problem by multiplying gate values to the expert outputs, adding a connection between the gating network and experts. For the scene representation, distinct scene regions are handled by different NeRF experts. The gating network routes 3D points to relevant experts based on gate values. A unified output head is designed to unify the prediction of multiple NeRF experts. This head imposes feature alignment among the implicit features of NeRF experts, resolving inconsistencies among different partitions.
Heterogeneous Mixture of Hash Experts (HMoHE) to model heterogeneous scene representations. Previous large-scale NeRF methods typically use the same network structure for different experts. However, different parts of the large-scale scenes have heterogeneous and diverse distributions, which need heterogeneous representations. In our framework, the learned decomposition of the entire scene is represented by a mixture of heterogeneous hash experts (HMoHE). Each decomposed scene part is handled by a distinct expert network. To ensure the diversity and flexibility of representation, we design the hash experts as heterogeneous. We make the experts exploit diverse resolution ranges, allowing our network to select a suitable resolution range for each scene part.
@@
Efficient implementation of hash-based gating and dispatching. To implement HMoHE more efficiently, we seamlessly integrate the dispatching of 3D points into the CUDA implementation of hash. We merge the hashing and dispatching of points together and avoid extra allocation of memory. This ensures that the point dispatching does not significantly impact the training efficiency.
@@
This paper significantly extends our previous ICLR version, i.e., Switch-NeRF [4]. Specifically, this work introduces a brand new design of the Heterogeneous Mixture of Hash framework. We redesign the gating network and experts in Switch-NeRF, exploiting highly efficient hash networks. We propose to learn heterogeneous decomposition and representation with the hash gating network and heterogeneous experts. For the point dispatching, we integrate it into the hash operation. We also adapt a new large-scale scene for better evaluation. Compared to Switch-NeRF, our new framework effectively models the heterogeneous representation of large-scale scenes and significantly improves the accuracy and efficiency. The proposed method achieves an 8x acceleration in training and a 16x acceleration in rendering compared to Switch-NeRF.
@@
  • We introduce a novel framework for large-scale neural radiance fields (NeRF), which integrates learnable scene decomposition, heterogeneous representation learning, and efficient modeling into a unified end-to-end architecture. Unlike prior methods, our framework does not rely on hand-crafted heuristic decomposition and is generally applicable to large-scale scene modeling.
@@
  • As the core of our framework, we design a hash-based gating network that learns to decompose scenes automatically and efficiently. We introduce Heterogeneous Mixture of Hash Experts (HMoHE) to enable learning scale-aware distributions. The gating and expert modules are jointly optimized to produce highly scalable and heterogeneous NeRF representations.
@@
  • We present an efficient implementation of the framework to make it a highly scalable and efficient learning framework that can handle extremely large-scale scenes. We adapt a new dataset for benchmarking highly scalable NeRF scene modeling. Our NeRF framework achieves significant improvement in terms of scalability, accuracy, and efficiency compared to the best-performing methods such as Switch-NeRF.
@@
Figure 2: Large-scale NeRFs with different decomposition methods. Dotted lines indicate non-differentiable operations, while solid lines represent differentiable operations that can back-propagate gradients in training. Mega-NeRF [2] divides the scene by clustering image pixels by 3D distances. Block-NeRF [3] divides image sets according to street blocks, which is actually based on camera distributions. In both kinds of methods, the decompositions are hand-crafted heuristics rules and not trainable. They require separate training for their sub-networks. Switch-NeRF introduces a unified framework where scene decomposition is learned automatically via a gating network, enabling end-to-end training of the entire framework. However, Switch-NeRF designs homogeneous experts with the same structures, which learns homogeneous representations of the scene. In contrast, Switch-NeRF++ employs an efficient hash network as the gate network and designs heterogeneous hash experts to learn heterogeneous decomposition and representation of a large-scale scene.
@@
Figure 3: The framework overview of Switch-NeRF and Switch-NeRF++ for large-scale scene modeling. In Switch-NeRF++, A 3D scene point x is first fed into the gating network to obtain a feature encoding $f_{g}$. $f_{g}$ is then sent to a small MLP head to predict gating values. The Top-11 index $i$ of the gate values is used in 3D point dispatching that determines which hash expert should be selected for the input point. In this way, the entire scene can be decomposed into different partitions. Then x is dispatched into its corresponding hash expert to encode its distinct distribution into a hash expert encoding $E_{i}(\textbf{x})$. Then we multiply the gate value $G_{i}(\textbf{x})$ on $E_{i}(\textbf{x})$ and obtain the final hash expert output $\hat{E}_{i}(\textbf{x})$, which is further fed into a small MLP head to predict the point density $\sigma$ and point color c. The spherical harmonics encoding (i.e., SH) of direction d, and the appearance embedding (i.e., AE) are used for view-dependent prediction of color c. In Switch-NeRF, the gating network and the experts are all MLP-based networks, which learns homogeneous representations. In contrast, in Switch-NeRF++, we leverage the efficient Hash-based Gating Network to dispatch 3D points. Furthermore, we design heterogeneous hash experts that have different scale ranges to explicitly model the heterogeneous distribution of the large-scale scene.
@@ 2 Related Work
Neural Radiance Fields. The Neural Radiance Fields method (NeRF) [1] is proposed to encode a 3D scene into an MLP network for novel view synthesis, with the multiple images and their cameras as input. The implicit framework of NeRF is very simple and effective, and has been widely applied to many 3D reconstruction tasks.
@@
Some NeRF variants apply NeRF to more challenging scenes. NeRF++ [28] extends NeRF to unbounded scenes by separating a scene into a foreground and a background, and applying the inverted sphere parameterization on the background. NeRF in the wild [29] attaches an appearance embedding to each image. It conditions the NeRF network with these appearance embeddings to model the appearance variation in each image in the wild environment. Mip-NeRF [30] and Mip-NeRF 360 [31] address sampling and aliasing of bounded and 360-degree scenes by designing a new scale-aware positional encoding. DSNeRF [32] uses sparse depth from colmap to regularize and speed up NeRF training. SPARF [33] and CVT-xRF [34] largely reduce the view number used in NeRF training and produce high-quality NeRF with sparse image inputs.
@@
Although the original NeRF targets novel view synthesis, it is a flexible 3D representation encoding 3D geometry. Several methods have focus on improving the geometry quality of NeRF for accurate 3D surface reconstruction [35, 36]. NeuS [35] represents a surface as a signed distance function (SDF) and learns an SDF field by volume rendering. Neuralangelo [36] recovers high-quality surface details by higher-order derivatives and coarse-to-fine optimization on the hash grids.
@@
NeRF methods typically need camera parameters as input. These cameras are usually predicted by SFM methods such as Colmap [37]. Several methods try to estimate camera parameters together with the neural radiance fields. NeRF– [38] treats camera parameters as trainable model parameters and jointly trains cameras in NeRF training. BARF [39] finds that using coarse-to-fine positional encoding in training helps the convergence of camera parameters.
@@
The training and inference speed of NeRF is critical in real-world applications. Many methods have been proposed for fast training and real-time rendering of NeRF. PlenOctrees [40], FastNeRF [41], and Kilo-NeRF [42] mainly speed up inference of NeRF. There are also methods to speed up both the training and rendering of NeRF, such as the DVGO [6], Plenoxels [43], Instant-NGP [5], TensoRF [8], and 3D Gaussian Splatting [44]. Instant-NGP uses a multi-resolution hash encoding to encode a scene into a set of hash maps by interpolating trainable hash features. TensoRF [8] represents a 3D radiance field using trainable 2D plane features and 1D feature vectors. Different from typical NeRF methods using volume rendering, 3D Gaussian Splatting (3DGS) [44] used 3D Gaussians to represent a 3D scene. It renders images by projection and $\alpha$-blending, which is very fast and supports high-resolution real-time rendering. 3DGS needs a point cloud from SfM or random initialization to initialize 3D Gaussians.
@@
The original NeRF needs to be trained on each scene and cannot generalize across different scenes. Pixel-NeRF [45] proposes to include image features as an additional condition and trains the model on a set of different scenes. MVSNeRF [46], NerfingMVS [47], and IBRNet [48] et. al. handle the unsupervised Multi-view Stereo problem by combining volume rendering and Multi-view Stereo together. These methods take single-view or multi-view image features as input and can generalize to different scenes.
@@
The flexible NeRF representation has been applied to 3D generation [49, 50, 51] and editing tasks [52, 53]. EG3D [49] simplifies the NeRF network into triplane feature maps, which have regular shapes and are easy for generation. LRM [50] proposes to use a large transformer to generate triplanes from a single image.
@@
The above-mentioned reconstruction methods typically work on small-scale scenes. Our method aims to scale NeRF methods to large-scale scenes and handle the problems of learnable decomposition, and heterogeneous, scalable, and efficient scene modeling for large-scale NeRF.
@@
Large-scale Neural Radiance Fields. When applied to large-scale scenes, a critical problem of NeRF is that handling a large data volume requires a scalable framework and more network parameters. Recent large-scale NeRF methods mainly focus on decomposing a large-scale scene into several parts and handling each part by a sub-network. Mega-NeRF [2] and Block-NeRF [3] successfully scale NeRF to urban-scale scenes sparsely by scene decomposition. Although achieving impressive accuracy, their scene decomposition is mainly conducted by heuristic hand-crafted rules. Mega-NeRF [2] places several centroids in a scene and divides the 3D points in the scene by 3D distances to these centroids. Block-NeRF [3] divides the training images into several subsets according to street blocks. These two methods train the sub-networks separately. Therefore, they need composition strategies such as distance-based weights to merge the rendering results. Different from methods using handcrafted decomposition methods, Switch-NeRF [4] firstly proposes a large-scale NeRF framework that automatically learns scene decomposition together with the Neural Radiance Fields in an end-to-end manner. It designs a sparse network structure containing a gating network and several NeRF experts, based on the mechanism of Mixture of Experts (MoE) [11]. It requires no prior knowledge of scene structures and decomposition rules. Our method extends from Switch-NeRF with brand-new designs of the hash-based network, heterogeneous experts, and efficient hash dispatching. The Mega-NeRF, Block-NeRF and our Switch-NeRF++ share a divide-and-conquer paradigm for large-scale scene reconstruction. However, we use fundamentally different decomposition methods and training strategies to scale the NeRF. As illustrated in Fig. 2, our Switch-NeRF++ introduces a unified framework that incorporates the decomposition and scene representation into a single learning network. It not only learns the scalable and heterogeneous scene representation by sparse experts, but also learns the heterogeneous decomposition itself. With the hash-based network, heterogeneous experts and efficient hash dispatching, and unified heads, our method demonstrates strong abilities to learn an accurate and consistent large-scale NeRF on various large-scale scenes. In contrast, Mega-NeRF and Block-NeRF use handcrafted rules to decompose the scene. Each of its sub-networks is trained independently. This makes the decomposition not aware of the scene structure, and the learned representations are not consistent across scenes. Our method remarkably improves the scalability, efficiency, and accuracy of large-scale NeRFs compared to best-performing methods.
@@
Some other methods [54, 9] also use plane features and hash encodings on large-scale scenes. However, they still encode a large-scale scene into a global network, so their scalability is still limited. The Gaussian splatting [55] has inspired large-scale NeRFs such as the Hierarchical-3D-Gaussians [56], CityGaussian [57], DoGaussian [58], VastGaussian [59], PyGS [60], OccluGaussian [61], etc. due to its efficiency. The Hierarchical-3D-Gaussians [56] uses a divide-and-conquer method that trains very large scenes in independent chunks. It organizes the chunks into a hierarchy and provides an efficient Level-of-Detail (LOD) rendering. CityGaussian [57] and VastGaussian [59] also use a divide-and-conquer method to manually divide the scene into smaller cells. In this paper, our method focuses on scaling to large-scale scenes by learning efficient scene decomposition and representations. Our method explores the better use of the multi-scale nature of grids to adaptively model the multi-scale nature of 3D scenes.
@@
Mixture of Experts (MoE). The methodology of Mixture of Experts has a long history for scaling the model size [62, 63]. Recent applications of MoE in deep learning mainly follow the design of [11]. It replaces the feed-forward networks in a language model with Mixture-of-Experts layers. The experts are activated sparsely by a Top-$k$ gating network. Each sample will only be dispatched to $k$ experts while $k$ is typically set as 2 and is much smaller than the total number of experts. In order to avoid imbalanced training and training collapse, it regularizes gate values by a gate loss, which can make the training of different experts more balanced. The MoE network is extensively applied for Natural Language Processing (NLP) [64, 12] and Vision [65, 66] tasks to scale the capacity and capability of neural networks. GShard [64] standardize the dispatching operations by Einops [67, 64]. It supports local dispatching and parallel gating by designing a simpler balancing loss and defining the capacity factor for experts. The value of $k$ in MoE gating directly influences the sparsity, efficiency, and accuracy of MoE. Top-$2$ typically can get good accuracy while maintaining sparsity [11, 64]. Switch Transformer [12] has investigated the training recipes of MoE in NLP tasks. It concludes that under a carefully designed training setting, Top-$2$ operation in MoE is not necessary. Top-$1$ in MoE can already get high accuracy while largely decreasing the cost of computation and communication for sample dispatching. Several MoE methods have investigated different gating mechanisms, such as Hash Routing [68], BASE [69], and Expert Choice Routing [70].